# III. Interacting Dark Energy: Summary of Models, Pathologies, and Constraints

### v1

@@ 3 Conditions to avoid pitfalls in the parameter space for each IDE model
For seven of the eight models, energy transfer from DM to DE results in negative densities. This creates a strong theoretical preference for scenarios where energy flows in the opposite direction, from DE to DM. Energy flow from DE to DM also alleviates the coincidence problem and makes big rip future singularities less likely.
@@
$3H(\delta_{\text{dm}}\rho_{\text{dm}}+\delta_{\text{de}}\rho_{\text{de}})$ DE $\rightarrow$ DM $\delta_{\text{dm}}>0;\;\delta_{\text{de}}>0;\;\delta_{\text{dm}}r_{0}+\delta_{\text{de}}<-\frac{wr_{0}}{(1+r_{0})}$
@@
$Q=3H(\delta_{\text{dm}}\rho_{\text{dm}}+\delta_{\text{de}}\rho_{\text{de}})$ $\delta_{\text{dm}}\left(3w+1\right)-\delta_{\text{de}}>w+\frac{1}{3}$ $\delta_{\text{dm}}\left(w+1\right)-\delta_{\text{de}}>w+1$
@@
$Q=3H\delta(\rho_{\text{dm}}-\rho_{\text{de}})$ $\delta>\frac{\frac{1}{3}+w}{2+3w}$ ; $\left[\delta\geq 0.666\right]$ $\delta>\frac{1+w}{2+w}$ ; $\left[\delta>-0.111\right]$
@@
Table 3: Conditions required to ensure accelerated expansion and to avoid a future big rip singularity for the different interaction kernels.
@@ 5 Conclusions
In Section 3, for all eight interactions, we provide: conditions to avoid imaginary or undefined energy densities (Table 1); conditions to avoid negative energy densities (Table 2); conditions to ensure accelerated expansion and avoid a future big rip (Table 3); and a summary of how these interactions address the coincidence problem (Table 4). From these results, we find that scenarios with energy flow from DE to DM are often theoretically more stable, as they tend to avoid negative energies and future singularities, alleviate the coincidence problem, and are broadly consistent with thermodynamic considerations [243]. Nevertheless, whether such scenarios are realized in nature must ultimately be decided by data, and future observational tests will be decisive.

### v3

@@ 3 Conditions to avoid pitfalls in the parameter space for each IDE model
The conditions in Table 1, 2 and 3 are illustrated in Figure 1, easing comparison between models and providing a more intuitive understanding of the parameter space of each model. For seven of the eight models, energy transfer from DM to DE ($\delta<0$) results in negative densities. This creates a strong theoretical preference for scenarios where energy flows in the opposite direction, from DE to DM ($\delta>0$). Energy flow from DE to DM also alleviates the coincidence problem and makes big rip future singularities less likely.
@@
$3H(\delta_{\text{dm}}\rho_{\text{dm}}+\delta_{\text{de}}\rho_{\text{de}})$ DE $\rightarrow$ DM $\delta_{\text{dm}}\geq 0;\;\delta_{\text{de}}\geq 0;\;\delta_{\text{dm}}r_{0}+\delta_{\text{de}}\leq-\frac{wr_{0}}{(1+r_{0})}$
@@
$Q=3H(\delta_{\text{dm}}\rho_{\text{dm}}+\delta_{\text{de}}\rho_{\text{de}})$ $\delta_{\text{dm}}\left(3w+1\right)-\delta_{\text{de}}\geq w+\frac{1}{3}$ $\delta_{\text{dm}}\left(w+1\right)-\delta_{\text{de}}\leq w+1$
@@
$Q=3H\delta(\rho_{\text{dm}}-\rho_{\text{de}})$ $\delta\geq\frac{\frac{1}{3}+w}{2+3w}$ ; $\left[\delta\geq 0.666\right]$ $\delta\leq\frac{1+w}{2+w}$ ; $\left[\delta\leq-0.111\right]$
@@
Table 3: Conditions required to ensure accelerated expansion and to avoid a future big rip singularity for the different interaction kernels. Values in square brackets are example values, where we have set $r_{0}=0.388$ and w=1w=-1.
@@
Figure 1: 2D Portraits of the parameter space for each of the 8 IDE models, using Table 1, 2 and 3. Blue areas indicate where the model has positive energy densities throughout all of cosmic evolution. Pink areas indicate that negative energies will occur in either the past or future expansion. The gray overlay indicates the presence of imaginary energy densities, while the purple areas show undefined energy densities. Lastly, the green mesh indicates the presence of future big rip singularities. The purple areas in the bottom two panels may have undefined values at some scale factor, but it is not guaranteed, see equations (2.23), (2.24), (2.26) and (2.27). For all cases $\Omega_{\rm{(dm,0)}}=0.266$, $\Omega_{\rm{(de,0)}}=0.685$ (implying $r_{0}=0.388$). For linear IDE model 1, we set w=1w=-1.
@@ 5 Conclusions
In Section 3, for all eight interactions, we provide: conditions to avoid imaginary or undefined energy densities (Table 1); conditions to avoid negative energy densities (Table 2); conditions to ensure accelerated expansion and avoid a future big rip (Table 3); and a summary of how these interactions address the coincidence problem (Table 4). These results are visualised in Figure 1. From these results, we find that scenarios with energy flow from DE to DM are often theoretically more stable, as they tend to avoid negative energies and future singularities, alleviate the coincidence problem, and are broadly consistent with thermodynamic considerations [247]. Nevertheless, whether such scenarios are realized in nature must ultimately be decided by data, and future observational tests will be decisive. For now, some discussions on the possibility of negative DE can be found in [248, 249, 250, 251, 37, 252].